# Multi-Modal Federated Learning for Cancer Staging over Non-IID Datasets with Unbalanced Modalities

### v1

@@ VI Conclusion and Future Works
In this work, we developed the system model for multi-modal FL over institutions with non-iid data and unbalanced number of modalities. To address disparate convergence rates across the various modalities involved in the system, we proposed DGB, weighing the gradients of different modalities in a non-uniform fashion. Further, to address the impact of non-iid data and the local model bias on the performance of DGB, we proposed PCW which takes into account the quality of data at institutions for refining the system-wide loss estimates fed to DGB. We demonstrated that our proposed methodology can outperform the state-of-the-art methods. Our proposed DGB and PCW methods are generic and can be used in other tasks such as Alzheimer’s disease prediction with data modalities of CT and PET scans, the exploration of which is left as future work. Other future directions include using generative adversarial networks (GANs) to replicate the missing modalities at each institution. Also, introducing distributed modality-aware attention mechanisms is a promising research direction.

### v2

@@ VI Conclusion and Future Works
In this work, we developed the system model for multi-modal FL over institutions with non-iid data and unbalanced number of modalities. To address disparate convergence rates across the various modalities involved in the system, we proposed DGB, weighing the gradients of different modalities in a non-uniform fashion. Further, to address the impact of non-iid data and the local model bias on the performance of DGB, we proposed PCW which takes into account the quality of data at institutions for refining the system-wide loss estimates fed to DGB. We demonstrated that our proposed methodology can outperform the state-of-the-art methods. Our proposed DGB and PCW methods are generic and can be used in other tasks such as Alzheimer’s disease prediction with data modalities of CT and PET scans, the exploration of which is left as future work. Another direction of research in this area is improving the feature extraction method used for each modality of data according to its distribution across the institution in the FL setting. Other future directions include using generative adversarial networks (GANs) to replicate the missing modalities at each institution are also worth looking into. Also, introducing distributed modality-aware attention mechanisms can be a promising research direction.